# On-Device Self-Supervised Learning of Low-Latency Monocular Depth from Only Events

### v1

@@ 4.1 Benchmarks
While our accuracy on the DSEC disparity benchmark falls short of supervised baselines, qualitative comparisons in Fig. 4 demonstrate that our approach effectively captures meaningful structures within disparity maps, even without ground truth labels during training. Notably, close objects—such as the car in interlaken_00_a(540) and traffic signs in thun_01_b(400), interlaken_01_a(1680), zurich_city_12_a(400) are accurately represented. Although the boundaries in our results may lack the sharpness achieved by supervised baselines, our approach better preserves contour shapes, such as the front of the car in interlaken_00_a(540), the arc of the tunnel in interlaken_00_b(560) and the pole in zurich_city_13_a(260). This advantage is especially evident for thin objects like the sign pole in interlaken_01_a(1680) and zurich_city_12_a(440), which are often challenging for supervised methods to capture accurately. Additionally, our self-supervised approach is immune to artifacts typically caused by the sporadic availability of ground truth at the image frame boundaries, resulting in smoother disparity maps free from discontinuity artifacts.
@@
Several factors constrain the accuracy of our methods on these benchmarks, including the reliance on self-supervised learning with events only, a compact network architecture, the use of monocular depth estimation rather than stereo and the imperfect estimation of scaling factor for converting monocular depth to metric depth. While our aim is not to surpass state-of-the-art methods, we believe the quality of our depth predictions is sufficient to support downstream tasks like robot navigation.
4.2 Robotic Experiments
@@
We first pretrain our network on the UZH-FPV dataset [9] using our self-supervised pipeline. This dataset was chosen for its diverse set of motion trajectories, enabling the network to learn a latent representation that generalizes well across various motion types. After pretraining, the network is deployed on a drone, where online learning (fine-tuning) is performed during flight. The network’s forward pass operates at an average speed of 30 Hz, with a backward pass and gradient update conducted every 10 forward passes. Following each gradient update, the state of the recurrent network is detached from the computational graph without resetting it. This approach ensures bounded memory usage, mitigates potential issues with gradient explosion or vanishing, and allows the network to retain temporal context effectively.
(a) Initialize from a network with random weights. Intervention rate: 50%.
(b) Initialize from pre-trained network, disable online learning. Intervention rate: 35%.
(c) Initialize from pre-trained network, enable online learning. Intervention rate: 11%.
Figure 5: Top-view flight trajectories from various experiments. Blue trajectories represent autonomous flights, while orange trajectories indicate manual flights where human intervention was necessary to prevent potential collisions. The setup with online learning demonstrates the best obstacle avoidance performance, achieving the lowest intervention rate.
Figure 6: Qualitative visualization of disparity map evolution during online learning. Note that the image is for visualization purposes only, as disparity estimation is event-based. The same colour map is applied to predictions from all different models for easy comparison. The pre-trained network begins at step 0, followed by 12K steps of online learning (OL) with streaming event data during flight. For comparison, we also show the prediction quality of a randomly initialized network trained from scratch (TFS) for 12K steps.
We show the metric improvements achieved during online learning in Fig. 7, where saved checkpoints are evaluated on a test sequence recorded in the same environment but with different placements for obstacles. The model shows significant improvement not only in the RSAT (ratio of squared average timestamps) metric, which is strongly correlated with the contrast maximization loss used to optimize the network, but also in MAE (mean absolute error) when compared against disparity ground truth. Additionally, the fine-tuning process is efficient, converging within just a few minutes of flight.
Qualitative results at different snapshots during online learning are presented in Fig. 6. Compared to step 0 (the pretrained model), the disparity values for certain close objects, such as the wall and poles, increase, as evidenced by the brightened colors in those regions. To highlight the advantages of pretraining, we also compare our model with a network initialized with random weights and trained using the same amount of online learning data, as shown in the last row of Fig. 6. The trained-from-scratch network fails to produce meaningful disparity maps within the flight’s limited timespan, underscoring the importance of pretraining for achieving fast and reliable adaptation during online learning.
@@
Figure 7: Mean absolute error (MAE) of disparity prediction and the ratio of squared average timestamp (RSAT) during online learning in flight. Model checkpoints were saved periodically and evaluated on a test sequence unseen by the model beforehand. 300 learning steps correspond to roughly 100 seconds of training during flight.
Finally, we plot the trajectories of the drone during the flight experiments under different settings in Fig. 5. During each experiment, the drone takes off and immediately starts to fly autonomously. A human pilot monitors the flight and intervenes if the drone is in a near collision with an obstacle. After the human pilot corrects for the collision course, the drone is switched to autonomous flying again. When the drone was initialized with random weights, the drone does not avoid obstacles and mostly fly with straight lines, as shown in Fig. 5(a). This resulted in constant human interventions and a highest intervention rate of 50 %, determined by the ratio of time of human piloting during a flight.
When we start the experiment with a pre-trained network, the obstacle avoidance performance improves dramatically, lowering the intervention rate to 35 % in Fig. 5(b). While the drone was already flying around obstacles, fine-tuning the model during flight further lowers the intervention rate to 11 %, as shown in Fig. 5(c), allowing the drone to smoothly fly between obstacles with more curvy trajectories than in previous other settings.
@@ 5 Conclusion
We have improved the efficiency of self-supervised learning of monocular depth estimation from camera events, such that on-device learning of low-latency monocular depth and ego-motion becomes feasible. The proposed approach features more efficient splatting, and has been implemented in CUDA instead of PyTorch. For common event rates (e.g., 10e6 Ev/s), it uses 100$\times$ less memory and computational time. When trained and benchmarked on event camera datasets, our small recurrent network outperforms other self-supervised approaches on the MVSEC dataset and captures essential structures with sufficient quality to support downstream navigation tasks on the DSEC dataset. Furthermore, we demonstrate online learning on-board a small flying drone, and show the usability of the resulting depth for obstacle avoidance. Finally, we investigate various combinations of pre-training and on-board learning, showing that the latter leads to a flight intervention rate reduction from 35% to 11%.
Our work scratches at the surface of the unused potential of on-board, online self-supervised learning. The current results already demonstrate that online learning leads to better performance in the operational environment. This confirms that learning with samples from the test environment will be important for challenging vision tasks such as monocular depth estimation from events. Moreover, it is important to note that the performance gap between self-supervised learning and supervised learning is still substantial (see Tab. 2). This suggests that there is ample room for improvement of the self-supervised learning scheme, which may lead to a larger performance improvement after online refinement. Moreover, improved detail in the estimated depth maps will facilitate more complex obstacle avoidance and navigation behaviors. Future work under such conditions would be even better positioned to investigate the downstream behavioral advantages of online self-supervised depth learning.
Self-supervised learning has turned out to be instrumental for learning advanced AI methods on enormous, unlabelled datasets. The current work is a small step towards unlocking this type of learning’s second major advantage, i.e., strongly reducing the need for generalization by learning in the test environment. We expect this to form a key component of truly autonomous robots that can quickly and aptly adapt to their environment.
@@
6 Optimized warping and splatting
Fig. 8 visualizes, on a high level, how our optimizations with respect to warping sets of events to different reference times are reducing runtime and memory consumption by preventing unnecessary warping operations.
(a) Warping events to multiple reference times in a batched manner involves zero-padding all bins to the same length (light-gray extensions), leading to unnecessary warping operations. Furthermore, events that were already warped out of the image space (and can be discarded) will still be warped.
(b) Warping all events individually in parallel does away with padding and allows to warp only those events that are still in the image space, giving the minimal amount of warping operations necessary.
Figure 8: Comparison of warping three bins of events to three reference times. Events originating from different bins have different colors. At every stage of warping (1, 2), events get warped to a neighboring bin, until we have reached all three reference times.
7 Extra DSEC qualitative results
We present additional qualitative results from the DSEC test set in Fig. 9. While the self-supervised results exhibit less sharp boundaries, they are free from many artifacts commonly introduced by supervised learning, such as discontinuities at frame boundaries and difficulties in handling thin objects.
8 More robot experiment results
Figs. 10, 11, 12 and 13 show extended qualitative results for four unseen scenes during a flight experiment. Looking at Fig. 10, we see that network’s ability to maximize the contrast of the image of warped events increases quickly from ‘step 0’ (only pre-training) to ‘step 3k’ (pre-training + 100 seconds of learning). While this improvement in terms of contrast maximization loss may be mostly due to just learning to increase the magnitude of the optical flow (as shown by the change from orange to purple in column 3, and black to red in column 5) through scaling depth and ego-motion, subsequent learning steps improve the depth map in more sophisticated ways, judging from the disappearance of the wrong ‘depth gap’ in the center of the disparity images in the third and fourth column. Similar priorities in learning (first scaling, then the rest) can be seen in the other scenes.
9 Training details
For offline training on datasets, we train for 50 epochs with the Adam optimizer and a learning rate of 1e-4. For contrast maximization, we accumulate 10 bins of events, and warp all events to all bin edges. Furthermore, we set the weight for the geometric consistency loss $\lambda=0.05$. For on-device learning, we lower the learning rate to 1e-5. Other specifics vary per dataset, and are mentioned below. In all cases, event streams are rectified.
@@ 10 Network details
Table 3: Network layer details. Unless specified otherwise, we use ELU activations, and a kernel size of 3, biases and ‘reflect’ padding for convolutional layers. Total parameter count is 430,416.
@@
Upsample(scale=8, ‘bilinear’) $(1,H/8,W/8)$ $(1,H,W)$
@@
11 Robot details
We built a 5-inch quadrotor for our robotic experiments. The drone is equipped with on-board sensors that provide the flight controller with all relevant information to follow high-level control commands. More specifically, the EKF running on the Kakute H7 Mini flight controller fuses IMU measurements with velocity and height measurements coming from a MTF-01 optical flow/range sensor into a stable position and velocity estimate. This allows a neural network (like our depth network) to give high-level commands like velocity setpoints or rotational rates.
All relevant components on the drone, along with their weight and power consumption, can be found in Tab. 4. All separate modules on the Jetson Orin NX are connected via ROS2, which also allows for logging to rosbags. ROS2 can also be connected via UART to the internal publish-subscribe messaging API that is used in the PX4 flight controller firmware.
@@ 11 Robot details
Table 4: List of hardware components used during robotic experiments. Power consumption estimates are obtained from Jetson’s jtop, battery drain during flight experiments, or component datasheets.
@@
Figure 10: Extended qualitative results on unseen data from a flight test recording. From top to bottom, we evaluate a pre-trained-only network (‘step 0’), then four networks after increasing amounts of online learning (OL), and finally a network trained-from-scratch. The bottom row starts with the single 20 ms bin of events currently seen by the network. The rest of the second column shows the accumulated events (multiple bins) in the current contrast maximization loss window. Applying the iterative warp by the optical flow constructed from depth and ego-motion gives the warped and deblurred events in the first column. Columns three and four show disparity at a global (colormap shared between rows) and local (colormap for only that row) scale. The last two columns show optical flow constructed from depth and ego-motion with global and local colormaps, respectively.
@@
Figure 11: Extended qualitative results on unseen data from a flight test recording. From top to bottom, we evaluate a pre-trained-only network (‘step 0’), then four networks after increasing amounts of online learning (OL), and finally a network trained-from-scratch. The bottom row starts with the single 20 ms bin of events currently seen by the network. The rest of the second column shows the accumulated events (multiple bins) in the current contrast maximization loss window. Applying the iterative warp by the optical flow constructed from depth and ego-motion gives the warped and deblurred events in the first column. Columns three and four show disparity at a global (colormap shared between rows) and local (colormap for only that row) scale. The last two columns show optical flow constructed from depth and ego-motion with global and local colormaps, respectively.
@@
Figure 12: Extended qualitative results on unseen data from a flight test recording. From top to bottom, we evaluate a pre-trained-only network (‘step 0’), then four networks after increasing amounts of online learning (OL), and finally a network trained-from-scratch. The bottom row starts with the single 20 ms bin of events currently seen by the network. The rest of the second column shows the accumulated events (multiple bins) in the current contrast maximization loss window. Applying the iterative warp by the optical flow constructed from depth and ego-motion gives the warped and deblurred events in the first column. Columns three and four show disparity at a global (colormap shared between rows) and local (colormap for only that row) scale. The last two columns show optical flow constructed from depth and ego-motion with global and local colormaps, respectively.
@@
Figure 13: Extended qualitative results on unseen data from a flight test recording. From top to bottom, we evaluate a pre-trained-only network (‘step 0’), then four networks after increasing amounts of online learning (OL), and finally a network trained-from-scratch. The bottom row starts with the single 20 ms bin of events currently seen by the network. The rest of the second column shows the accumulated events (multiple bins) in the current contrast maximization loss window. Applying the iterative warp by the optical flow constructed from depth and ego-motion gives the warped and deblurred events in the first column. Columns three and four show disparity at a global (colormap shared between rows) and local (colormap for only that row) scale. The last two columns show optical flow constructed from depth and ego-motion with global and local colormaps, respectively.

### v2

@@ Results.
While our accuracy on the DSEC disparity benchmark falls short of supervised baselines, qualitative comparisons in Fig. 4 demonstrate that our approach effectively captures meaningful structures within disparity maps, even without ground truth labels during training. Notably, close objects, such as the car in interlaken_00_a(540) and traffic signs in thun_01_b(400), interlaken_01_a(1680) and zurich_city_12_a(400), are accurately represented. Although the boundaries in our results may lack the sharpness achieved by supervised baselines, our approach better preserves contour shapes, such as the front of the car in interlaken_00_a(540), the arc of the tunnel in interlaken_00_b(560) and the pole in zurich_city_13_a(260). This advantage is especially evident for thin objects like the sign pole in interlaken_01_a(1680) and zurich_city_12_a(440), which are often challenging for supervised methods to capture accurately. Additionally, our self-supervised approach can run at higher-than-ground-truth frequencies (100 Hz vs 10 Hz) and is immune to artifacts typically caused by the sporadic availability of ground truth at the image boundaries, resulting in smoother disparity maps free from discontinuity artifacts.
Limitations.
@@
Several factors constrain the accuracy of our methods on these benchmarks, including the reliance on self-supervised learning with events only, a compact network architecture, the use of monocular depth estimation rather than stereo and the imperfect estimation of a scaling factor for converting monocular depth to metric depth. Errors in few-event areas could be reduced by, e.g., including the reconstruction loss from [27]. However, our aim is not to surpass state-of-the-art methods, and we believe the quality of our depth predictions is sufficient to support downstream tasks like robot navigation.
4.2 Drone experiments
Setup.
@@ 4.2 Drone experiments
We first pre-train our network on the UZH-FPV dataset [9] using our self-supervised pipeline. This dataset was chosen for its diverse set of motion trajectories, enabling the network to learn a latent representation that generalizes well across various motion types. After pre-training, the network is deployed on a drone, where online learning (fine-tuning) is performed during flight. The network’s forward pass operates at an average speed of 30 Hz, with a backward pass and gradient update conducted every 10 forward passes. During flight experiments, we set the drone to fly at a constant height and a forward speed of 0.5 m/s. The predicted depth is binned and used to control the drone’s yaw rate.
@@
Figure 5: Left: Boxplots of distance between pilot interventions during flight experiments. While using ground truth (GT) depth is best, adding online learning (PT + OL) improves over just pre-training (PT) by ~30%. Training from scratch (TFS) does not result in meaningful obstacle avoidance. Right: MAE (mean absolute error) of depth prediction and RSAT (ratio of squared average timestamps, indicates deblurring quality) during online learning in flight. Model checkpoints were saved periodically and evaluated on a test sequence unseen by the model beforehand. 300 learning steps correspond to roughly 100 seconds of training during flight.
Results.
We show the quantitative improvements achieved through online learning in Fig. 5 (right plot), where saved checkpoints are evaluated on a test sequence recorded in the same environment but with different placements for obstacles. The model shows significant improvement not only in the RSAT (ratio of squared average timestamps) metric [19], which is strongly correlated with the contrast maximization loss used to optimize the network, but also in MAE (mean absolute error) when compared against ground truth depth. Additionally, the fine-tuning process is efficient, converging within just two minutes of flight.
Figure 6: Qualitative visualization of disparity map evolution during online learning. Note that the image is for visualization purposes only, as disparity estimation is event-based. The same color map is applied to predictions from all different models for easy comparison. The pre-trained network begins at step 0, followed by 12K steps of online learning (OL) with streaming event data during flight. For comparison, we also show the prediction quality of a randomly initialized network trained from scratch (TFS) for 12K steps.
Qualitative results at different snapshots during online learning are presented in Fig. 6. Compared to step 0 (the pre-trained model), the disparity values for certain close objects, such as the wall and poles, increase, as evidenced by the brightened colors in those regions. To highlight the benefits of pre-training, we also compare our model with a network initialized with random weights and trained using the same amount of online learning data, as shown in the last row of Fig. 6. The from-scratch network fails to produce meaningful disparity maps within the flight’s limited timespan, underscoring the importance of pre-training in achieving fast and reliable adaptation during online learning.
Finally, we quantify that these improvements in depth estimation through online learning translate to better obstacle avoidance performance in flight experiments. During each experiment, the drone takes off and immediately starts to fly autonomously. A human pilot monitors the flight and intervenes if the drone is in a near collision with an obstacle. After the human pilot corrects for the collision course, the drone is switched to autonomous flying again.
We compare the distance between pilot interventions for the different experiments in Fig. 5 (left plot), and show top views of the flight trajectories in Fig. 7. When training from scratch (network initialized with random weights), the drone does not avoid obstacles and mostly flies in straight lines. When we start with a pre-trained network, we see actual obstacle-avoiding behavior, and the distance between pilot interventions goes up by ~65%. When adding online learning during flight, the distance between interventions improves by a further ~30%, and the flight behavior seems to become more diverse.
Limitations.
The quality of the on-board depth maps is limited by the fact that only a quarter of the camera’s 640x480 resolution is used (compare Fig. 4 and Fig. 6). We mitigate this with an artificially textured environment to ensure sufficient motion-induced events. Higher-quality depth maps, or operation in more natural environments, will require using more of the camera’s resolution.
To further enhance computational efficiency and performance, the nonlinear motion model [28] could be traded for the cheaper-to-compute linear variant [19], at the cost of increased errors on nonlinear event trajectories. Furthermore, inference and learning could be run asynchronously at different rates [36], or only limited to partial network fine-tuning for a few selected layers. Incorporating a jointly optimized flow decoder tail [42] could also improve depth estimation for dynamic obstacles.
Dynamic objects moving towards the drone would already be avoided by our current pipeline (even though they are not included in training). However, due to the static scene assumption, their depth is underestimated (like the oncoming van in Fig. 4).
Lastly, The current depth-based yaw control is attracted by corners in the environment, requiring pilot intervention (see the bottom left environment corners in Fig. 7). Solving this, or allowing for more complex environments (e.g., higher obstacle density), would require a more advanced control strategy capable of better interpreting depth cues.
Figure 7: Top-view flight trajectories of various experiments. Blue represents autonomous flight, while orange indicates pilot interventions necessary to prevent collisions/going out of bounds. Training from scratch does not give meaningful obstacle-avoiding behavior. Online learning results in longer autonomous sections and more diverse paths than just pre-training. Using ground truth depth (from RealSense) results in almost-perfect avoidance, and only requires intervention when flying into a corner (limitation of control algorithm).
@@ 5 Conclusion
We have improved the efficiency of self-supervised learning of monocular depth estimation from events, such that on-device learning of low-latency monocular depth and ego-motion becomes feasible. The proposed approach features more efficient and parallel processing, and has been implemented in CUDA instead of PyTorch. For common event rates (0.1-1 Mev/s) this reduces runtime by 100x, while using 2-5x less memory—improvements that would also transfer to other pipelines involving warping/splatting of events [32, 10, 27] and images [26].
When trained and benchmarked on event camera datasets, our small recurrent network outperforms other self-supervised approaches and captures essential structures with sufficient quality to support downstream navigation tasks. Furthermore, we demonstrate that online learning on board a small flying drone leads to improved depth estimates within two minutes of learning, leading to more successful obstacle avoidance (~30% improvement in distance between pilot interventions).
Our work taps into the unused potential of on-board, online self-supervised learning. The current results already demonstrate that online learning leads to better performance in the operational environment. While SSL still needs further improvements to match supervised baselines, its core advantages—pretraining on large unlabeled datasets and finetuning directly in the test environment—hold the key to truly robust autonomous robot deployment across diverse real-world settings.
Acknowledgments.
The authors would like to thank the reviewers for their constructive feedback and suggestions. This work was supported by funding from NWO (NWA.1292.19.298), the Air Force Office of Scientific Research (award no. FA8655-20-1-7044) and the Office of Naval Research Global (award no. N629092112014).
@@ Acknowledgments.
https://mavlab.tudelft.nl/depth_from_events
6 Extra qualitative results
DSEC.
We present additional qualitative results from the DSEC test set in Fig. 8. While the self-supervised results exhibit less sharp boundaries, they are free from artifacts commonly introduced by supervised learning, such as discontinuities at image borders and around thin objects.
Robot experiments.
Figs. 9, 10, 11 and 12 show extended qualitative results for four unseen scenes from flight experiments. Looking at Fig. 9, we see that network’s ability to maximize the contrast of the image of warped events increases quickly from “step 0” (only pre-training) to “step 3k” (pre-training + 100 seconds of learning). While this improvement in terms of contrast maximization loss may be mostly due to just learning the correct magnitude of the optical flow (as shown by the change from orange to purple in column 3, and black to red in column 5) through scaling depth and ego-motion, subsequent learning steps improve the depth map in more sophisticated ways, judging from the disappearance of the wrong “depth gap” in the center of the disparity images in the third and fourth column. Similar priorities in learning patterns can be seen in the other scenes.
7 Implementation details
Training.
For offline training on datasets, we train for 50 epochs with the Adam optimizer and a learning rate of 1e-4. For contrast maximization, we accumulate 10 bins of events, and warp all events to all bin edges. Furthermore, we set the weight for the geometric consistency loss $\lambda=0.05$. For on-device learning, we lower the learning rate to 1e-5. Specifics per dataset are mentioned below. In all cases, event streams are undistorted and rectified.
@@ 7 Implementation details
Upsample(scale=8, "bilinear") $(1,H/8,W/8)$ $(1,H,W)$
@@
Table 3: Network layer details. Unless specified otherwise, we use ELU activations, and a kernel size of 3, biases and “reflect” padding for convolutional layers. Total parameter count is 430,416.
Depth-based control.
We slice depth maps into $K=8$ vertical bins, compute a vector of average inverse depths (disparities) $\bm{d}\in\mathbb{R}^{8}$, and use it to set a target yaw rate. Furthermore, $\lambda_{\text{goal}}=0.2$, $\lambda_{\text{avoid}}=1.0$, $\alpha=0.5$, $\sigma=12.0$.
Drone setup.
We built a 5-inch quadrotor for our robot experiments. The drone is equipped with on-board sensors that provide the flight controller with all relevant information to follow high-level control commands. More specifically, the EKF running on the Kakute H7 Mini flight controller fuses IMU measurements with velocity and height measurements coming from an MTF-01 optical flow/range sensor into a stable position and velocity estimate. This allows a neural network (like our depth network) to give high-level commands like velocity setpoints or rotational rates.
All relevant components on the drone, along with their weight and power consumption, can be found in Tab. 4. Communication on the Orin is handled with ROS2 [25], which also allows for logging to rosbags, and can be connected via UART to the internal publish-subscribe messaging API of the PX4 flight controller firmware.
@@
Table 4: List of hardware components used during robot experiments. Power consumption estimates are obtained from Jetson’s jtop, battery drain during flight experiments, or component datasheets.
@@
Figure 9: Extended qualitative results on unseen data from a flight test recording. From top to bottom, we evaluate a pre-trained-only network (“step 0”), then four networks after increasing amounts of online learning (OL), and finally a network trained-from-scratch. The bottom row starts with the single 20 ms bin of events currently seen by the network. The rest of the second column shows the accumulated events (multiple bins) in the current contrast maximization loss window. Applying the iterative warp by the optical flow constructed from depth and ego-motion gives the warped and deblurred events in the first column. Columns three and four show disparity at a global (color map shared between rows) and local (color map for only that row) scale. The last two columns show optical flow constructed from depth and ego-motion with global and local color maps, respectively.
@@
Figure 10: Extended qualitative results on unseen data from a flight test recording. From top to bottom, we evaluate a pre-trained-only network (“step 0”), then four networks after increasing amounts of online learning (OL), and finally a network trained-from-scratch. The bottom row starts with the single 20 ms bin of events currently seen by the network. The rest of the second column shows the accumulated events (multiple bins) in the current contrast maximization loss window. Applying the iterative warp by the optical flow constructed from depth and ego-motion gives the warped and deblurred events in the first column. Columns three and four show disparity at a global (color map shared between rows) and local (color map for only that row) scale. The last two columns show optical flow constructed from depth and ego-motion with global and local color maps, respectively.
@@
Figure 11: Extended qualitative results on unseen data from a flight test recording. From top to bottom, we evaluate a pre-trained-only network (“step 0”), then four networks after increasing amounts of online learning (OL), and finally a network trained-from-scratch. The bottom row starts with the single 20 ms bin of events currently seen by the network. The rest of the second column shows the accumulated events (multiple bins) in the current contrast maximization loss window. Applying the iterative warp by the optical flow constructed from depth and ego-motion gives the warped and deblurred events in the first column. Columns three and four show disparity at a global (color map shared between rows) and local (color map for only that row) scale. The last two columns show optical flow constructed from depth and ego-motion with global and local color maps, respectively.
@@
Figure 12: Extended qualitative results on unseen data from a flight test recording. From top to bottom, we evaluate a pre-trained-only network (“step 0”), then four networks after increasing amounts of online learning (OL), and finally a network trained-from-scratch. The bottom row starts with the single 20 ms bin of events currently seen by the network. The rest of the second column shows the accumulated events (multiple bins) in the current contrast maximization loss window. Applying the iterative warp by the optical flow constructed from depth and ego-motion gives the warped and deblurred events in the first column. Columns three and four show disparity at a global (color map shared between rows) and local (color map for only that row) scale. The last two columns show optical flow constructed from depth and ego-motion with global and local color maps, respectively.